Commute times in the U.S. are heavily skewed to the right. We select a random sample of 220 people from the 2000 U.S. Census who reported a non-zero commute time.

In this sample the mean commute time is 28.3 minutes with a standard deviation of 19.2 minutes. Can we conclude from this data that the mean commute time in the U.S. is less than half an hour? Conduct a hypothesis test at the 5% level of significance.

What is the p-value for this hypothesis test?

Your answer should be rounded to 4 decimal places.
Answer #1

Solution :

Given that,

Population mean = \mu = 30

Sample mean = \bar x = 28.3

Sample standard deviation = s = 19.2

Sample size = n = 220

Level of significance = \alpha = 0.05

This is a left (One) tailed test,

The null and alternative hypothesis is,  

Ho: \mu \geq 30

Ha: \mu < 30

The test statistics,

t = ( \bar x - \mu )/ (s/\sqrt{n})

= ( 28.8 - 30 ) / ( 19.2 / \sqrt{} 220 )

= -1.313

df = n - 1 = 219

P- Value = 0.0952

The p-value is p = 0.0952 > 0.05, it is concluded that the null hypothesis is fail to rejected.